# Doppler Correspondence: Non-Iterative Scan Matching With Doppler Velocity-Based Correspondence

### v2

@@ IV-F Correspondence Visualization
In the sparse 4D radar point clouds, Doppler Correspondence generates significantly fewer correspondences than a number of points, as illustrated in Figs. 8(a) and (d). However, those correspondences are high quality, leading to better alignment performance compared with the closest-point method, as shown in Fig. 6(a). This result aligns with the less is more principle discussed in [8], where using fewer but more reliable correspondences can be advantageous in highly noisy data. For semi-dense 4D radar point cloud, while the closest point correspondence often fails to establish reliable matches, Doppler Correspondence succeeds in maintaining robust and consistent pairings, even when only a single iteration is performed, as illustrated in Figs. 8(b) and (e). For dense point clouds with repetitive geometric structures (Baker-Barry Tunnel (Empty)), Doppler Correspondence can produce clear and stable matches by effectively mitigating the ambiguity that may degrade the performance of ICP approaches, as depicted in Figs. 8(c) and (f).
@@ V-A Key Assumptions and Their Breakdown
$=-\Delta T\omega^{\times}q_{j}-(I-\Delta T\omega^{\times})\Delta t.$
@@
$=-\rho\cdot n_{q_{j}}+(\Delta T\omega^{\times}\rho)\cdot n_{q_{j}}$
$=-v_{q,j}+(\Delta T\omega^{\times}\rho)\cdot n_{q_{j}}.$
@@ VII Conclusion
TABLE II: Odometry performance on the semi-dense 4D radar dataset. Each cell reports the absolute pose errors [5] (in meters). Best results are in bold, second-best are underlined.
@@
Generalized-ICP [17] 64.21 57.28 138.41 3.30
VGICP [9] 50.86 61.86 141.37 4.09
DICP [7] 60.15 68.60 143.24 2.94
@@
TABLE III: Odometry performance on the dense 4D LiDAR dataset. Each cell reports the absolute pose errors [5] (in meters). Best results are in bold, second-best are underlined. Note that Div. indicates the trajectory of the algorithm diverged.
@@
Generalized-ICP [17] Div. Div.
VGICP [9] Div. Div.
DICP [7] 65.27 37.12

### v3

@@ IV-F Correspondence Visualization
In the sparse 4D radar point clouds, Doppler Correspondence generates significantly fewer correspondences than a number of points, as illustrated in Figs. 8(a) and (d). However, those correspondences are high quality, leading to better alignment performance compared with the closest-point method, as shown in Fig. 6(a). This result aligns with the less is more principle discussed in [huang2024less], where using fewer but more reliable correspondences can be advantageous in highly noisy data. For semi-dense 4D radar point cloud, while the closest point correspondence often fails to establish reliable matches, Doppler Correspondence succeeds in maintaining robust and consistent pairings, even when only a single iteration is performed, as illustrated in Figs. 8(b) and (e). For dense point clouds with repetitive geometric structures (Baker-Barry Tunnel (Empty)), Doppler Correspondence can produce clear and stable matches by effectively mitigating the ambiguity that may degrade the performance of ICP approaches, as depicted in Figs. 8(c) and (f).
@@ V-A Key Assumptions and Their Breakdown
$\Delta T\omega^{\times}q_{j}-(I-\Delta T\omega^{\times})\Delta t.$
@@
$\rho\cdot n_{q_{j}}+(\Delta T\omega^{\times}\rho)\cdot n_{q_{j}}$
$v_{q,j}+(\Delta T\omega^{\times}\rho)\cdot n_{q_{j}}.$
@@ VII Conclusion
TABLE II: Odometry performance on the semi-dense 4D radar dataset. Each cell reports the absolute pose errors [grupp2017evo] (in meters). Best results are in bold, second-best are underlined.
@@
Generalized-ICP [segal2009generalized] 64.21 57.28 138.41 3.30
VGICP [koide2021voxelized] 50.86 61.86 141.37 4.09
DICP [hexsel2022dicp] 60.15 68.60 143.24 2.94
@@
TABLE III: Odometry performance on the dense 4D LiDAR dataset. Each cell reports the absolute pose errors [grupp2017evo] (in meters). Best results are in bold, second-best are underlined. Note that Div. indicates the trajectory of the algorithm diverged.
@@
Generalized-ICP [segal2009generalized] Div. Div.
VGICP [koide2021voxelized] Div. Div.
DICP [hexsel2022dicp] 65.27 37.12